OpenStudy (lulubj):

Find the product. Write your answer in exponential form. 7^-1 x 7^-7

OpenStudy (welshfella):

when you multiply numbers with exponents you ADD the exponents For example 3^3 * 3^-1 = 3 ^(3+ (-1) = 3^2

OpenStudy (welshfella):

so in your case we have 7^-1 * 7^-7 = 7^(-1 + (-7))

OpenStudy (lulubj):

ok

OpenStudy (lulubj):

so what would it be

OpenStudy (welshfella):

well what is -1 + (-7) ?

OpenStudy (welshfella):

note that +( - = -

OpenStudy (supersmart1001):

WELCOME TO OPENSTUDY @lulubj

OpenStudy (welshfella):

yes welcome @lulubj we like to guide to the answer of your problem rather than just give you the answer. That way you learn something.

OpenStudy (welshfella):

-1 + (-7) = -1 - 7 = ?

OpenStudy (welshfella):

perhaps a number line will help you with this 

OpenStudy (welshfella):

what would be the value of x?

OpenStudy (welshfella):

Going down the line 1 unit corresponds to the -1.

OpenStudy (welshfella):

what is the next negative whole number below -7 ?

OpenStudy (lulubj):

-8?

OpenStudy (welshfella):

YES -7-1 = -8 so 7^-1 * 7*-7 = 7 ^-8 note that the 'base' ( that is number 7) stays the same.

OpenStudy (lulubj):

Thanks a lot

OpenStudy (welshfella):

another example for you is 4^-4 * 4^-2 = 4^(-4 + (-2) = 4^(-4-2) = 4^-6

OpenStudy (welshfella):

yw